with the mean of 0.503 and standard deviation of 0.004 what is the diameter that is exceeded by 53% of ball bearing
1
Expert's answer
2020-11-02T14:50:33-0500

P(x>X)=1-P(x<X)=0.53

P(x<X)=0.47

z-value for 0.47 is -0.07

(X-mean)/sd=-0.07

X=-0.07*0.004+0.503=0.5027